# Benchmarking Filtered Approximate Nearest Neighbor Search Algorithms on Transformer-based Embedding Vectors

### v1

@@ Abstract.
Advances in embedding models for text, image, audio, and video drive progress across multiple domains, including retrieval-augmented generation, recommendation systems, vehicle/person re-identification, and face recognition. Many applications in these domains require an efficient method to retrieve items that are close to a given query in the embedding space while satisfying a filter condition based on the item’s attributes, a problem known as Filtered Approximate Nearest Neighbor Search (FANNS). In this work, we present a comprehensive survey and taxonomy of FANNS methods and analyze how they are benchmarked in the literature. By doing so, we identify a key challenge in the current FANNS landscape: the lack of diverse and realistic datasets, particularly ones derived from the latest transformer-based text embedding models. To address this, we introduce a novel dataset consisting of embedding vectors for the abstracts of over 2.7 million research articles from the arXiv repository, accompanied by 11 real-world attributes such as authors and categories. We benchmark a wide range of FANNS methods on our novel dataset and find that each method has distinct strengths and limitations; no single approach performs best across all scenarios. ACORN, for example, supports various filter types and performs reliably across dataset scales but is often outperformed by more specialized methods. SeRF shows excellent performance for range filtering on ordered attributes but cannot handle categorical attributes. Filtered-DiskANN and UNG excel on the medium-scale dataset but fail on the large-scale dataset, highlighting the challenge posed by transformer-based embeddings, which are often more than an order of magnitude larger than earlier embeddings. We conclude that no universally best method exists.
Code:
https://github.com/spcl/fanns-benchmark
Datasets:
https://huggingface.co/datasets/SPCL/arxiv-for-fanns-small https://huggingface.co/datasets/SPCL/arxiv-for-fanns-medium https://huggingface.co/datasets/SPCL/arxiv-for-fanns-large
@@ 1. Introduction
Advances in embedding models for text (Lee et al., 2024; Lei et al., 2025; Zhang et al., 2024), image (Szegedy et al., 2015; Liu et al., 2024), video (Sivic and Zisserman, 2003a), audio (Cramer et al., 2019; Aytar et al., 2016), and other modalities (Lerer et al., 2019) have significantly enhanced semantic search, where items are mapped to a high-dimensional vector space, and similarity is measured by the distance between embedding vectors. Efficient similarity search algorithms are essential to navigate this space. Given the scale of modern datasets, exact nearest neighbor search (ENNS) algorithms are too slow, necessitating the use of approximate nearest neighbor search (ANNS) methods (Indyk and Motwani, 1998; Malkov et al., 2014; Malkov and Yashunin, 2018; Jayaram Subramanya et al., 2019; Sivic and Zisserman, 2003b; Jegou et al., 2010; Gong et al., 2020; Babenko and Lempitsky, 2014a; Guo et al., 2016; Ge et al., 2013).
@@
Many real-world applications, including large language models (LLMs) with retrieval-augmented generation (Gao et al., 2023; Xu et al., 2021; Shapkin et al., 2023), recommendation systems (Suchal and Návrat, 2010; Yang et al., 2020b; Koenigstein et al., 2012), vehicle and person re-identification (Liu et al., 2016a; Zuo and Deng, 2023), face and voice recognition (Grother et al., 2019; Yang et al., 2020b; Richardson et al., 2015), and e-commerce (Li et al., 2018; Wei et al., 2020), require retrieving only items that satisfy specific filtering conditions on item attributes, e.g., access rights for a document, a video’s timestamp, or a product’s price. These requirements have driven the development of filtered approximate nearest neighbor search (FANNS). Both academia and industry have recognized its growing importance, leading to numerous research publications (Wang et al., 2024; Mohoney et al., 2023; Gollapudi et al., 2023; Gupta et al., 2023; Zuo and Deng, 2023; Patel et al., 2024; Engels et al., 2024; Zuo et al., 2024; Xu et al., 2024; Liang et al., 2024) and adoption in database solutions such as Pinecone (Pinecone, 2025), Vespa (Vespa, 2025), and Vectara (Vectara, 2025). Evaluating FANNS methods requires vector datasets that reflect real-world workloads. Yet, despite the rapid rise of transformer-based text embeddings, no FANNS benchmarks with such vectors are known to us.
Our contributions are threefold. First, to clarify the requirements for effective FANNS benchmarks, we survey the current FANNS landscape and propose a comprehensive taxonomy of existing methods. We classify approaches along three key dimensions: filtering approach, indexing technique, and supported filter types. Second, we present a novel dataset featuring vectors generated by embedding paper abstracts from the arXiv dataset (Cornell University et al., 2025) using the transformer-based model stella_en_400M_v5 (NovaSearch, 2025; Zhang et al., 2024). Each item in our dataset has 111111 real-world attributes. Third, we benchmark a diverse set of FANNS methods on this dataset. Our evaluation reveals that each method involves distinct trade-offs, and no single approach dominates across all conditions. For example, ACORN supports multiple filter types and scales well, but is often outperformed by more specialized methods. Filtered-DiskANN and UNG perform strongly at medium scale but fail on our largest dataset. SeRF performs consistently well but supports only range filters on ordered attributes. These results demonstrate that FANNS remains a nuanced design space, and that selecting the right method depends on the specific workload and filtering requirements.
@@ 2. Background
2.1. FANNS Problem Statement
In FANNS, we are given $n$ items, $m\geq 1$ attributes, and $p$ queries. Each item $I_{i}=(v_{i},a_{i,1},...,a_{i,m})$ for $i\in\{1,\dots,n\}$ consists of a $d$-dimensional embedding vector $v_{i}$ and a value for each of the $m$ attributes ($a_{i,1}$ to $a_{i,m}$). An item can, e.g., represent a research paper with $v_{j}$ being an embedding of its abstract and attributes corresponding to venue, year of publication, and authors. A query $Q_{j}=(q_{j},k_{j},f_{j})$ for $j\in\{1,\dots,p\}$ consists of a $d$-dimensional query vector $q_{j}$, the number of items to return $k_{j}$, and a filter function $f_{j}$ that determines whether an item matches the filter. For a query, $q_{j}$ can, e.g., be the embedding of a search term with $f_{j}$ only returning true for papers submitted to VLDB. We use $o_{j}\leq m$ to denote the number of attributes the filter $f_{j}$ depends on. Table 1 summarizes all parameters.
Table 1. 2.1) Parameters used in FANNS.
Parameter Description
$n$ Number of items
$m$ Number of attributes
$p$ Number of queries
$d$ Dimensionality of embedding and query vectors
$v_{i}$ Embedding vector of the $i$-th item
$a_{i,l}$ $l$-th attribute of the $i$-th item
$q_{j}$ Query vector of the $j$-th query
$k_{j}$ Number of items to return for the $j$-th query
$f_{j}$ Filter expression of the $j$-th query
$o_{j}$ Number of attributes considered in $f_{j}$
To answer a query $Q_{j}$, we approximately retrieve the $k_{j}$ items that satisfy the filter $f_{j}$ and have embedding vectors closest to the query vector $q_{j}$ according to a given distance function (e.g., Euclidean distance, cosine similarity, Manhattan distance, or Hamming distance). Most FANNS algorithms return only items that match the filter; however, some approximate the filtering step, occasionally retrieving items that do not satisfy it. We refer to this relaxed version of the FANNS problem as the approximately filtered approximate nearest neighbor search (AFANNS) problem.
2.2. Example Use-Case of FANNS
To illustrate the practical use of FANNS, we introduce an example application depicted in Figure 1. We reference components and actions in Figure 1 using letters [Uncaptioned image] and numbers [Uncaptioned image], respectively. Consider a semantic search engine for research papers; a system that retrieves the most relevant papers for a given query, even when different terminology is used to describe the same concept. To provide fine-grained control over search results, we incorporate filters for venue, publication year, and authors.
@@ 2.2. Example Use-Case of FANNS
Index Construction: To build such a system, we transform each paper in our database [Uncaptioned image] into an item, as described in Section 2.1. The paper’s venue, publication year, and authors define [Uncaptioned image] the item’s attributes. The paper’s abstract is processed [Uncaptioned image] by a text embedding model [Uncaptioned image], such as NV-Embed (Lee et al., 2024; NVIDIA, 2025), LENS (Lei et al., 2025; Lei, 2025), Stella (Zhang et al., 2024; NovaSearch, 2025), GTE (Li et al., 2023; Alibaba-NLP, 2025), or BGE (Chen et al., 2024; of Artificial Intelligence, 2025), to generate the item’s embedding vector. To efficiently answer FANNS queries for these items, we insert [Uncaptioned image] them into a FANNS index [Uncaptioned image].
@@
Query execution: When a user enters [Uncaptioned image] a search query, we transform it into a FANNS query, as described in Section 2.1. The filter $f_{j}$ and the number of requested results $k_{j}$ are set [Uncaptioned image] based on the user’s input. The query text is processed [Uncaptioned image] by the same text embedding model [Uncaptioned image] used during index construction to generate the query vector. The query is submitted [Uncaptioned image] to a FANNS algorithm [Uncaptioned image], which utilizes [Uncaptioned image] the FANNS index [Uncaptioned image] to retrieve the IDs of the most relevant papers. These IDs are forwarded [Uncaptioned image] to a retriever [Uncaptioned image], which fetches [Uncaptioned image] the corresponding papers from the database [Uncaptioned image] and returns [Uncaptioned image] them to the user.
Figure 1. 2.2) Example use case of FANNS.
2.3. Attribute Types
We provide a list of attribute types supported by FANNS algorithms.
  • Unordered attributes: These attributes contain a single value from a domain with no total order. An example is the venue of a published paper, as each paper appears in a single venue, and no total order exists between venues.
  • Ordered attributes: These attributes contain a single value from a totally ordered domain. An example is the publication year of a paper, as each paper has a single publication year, and years follow a total order.
  • Set attributes: These attributes contain a set of values from an unordered domain. For instance, the authors of a research paper can be represented as a set attribute.
2.4. Filter Types
We categorize filters that apply to a single attribute into three types. Table 2 shows the compatibility of filters types with attribute types.
  • Exact match (EM) filter: Matches only if an item’s attribute value is equal to the queried value.
  • Range (R) filter: Used on an ordered attribute and matches only if an item’s attribute value is within the queried range.
  • Exact match in set (EMIS) filter: Applies to a set attribute and matches only if the item’s set attribute contains the queried value.
FANNS algorithms that apply filtering to multiple attributes of the same type implement multiple exact match (MEM), multiple range (MR), and multiple exact match in set (MEMIS) filters. We define a combined (C) filter as a filter capable of combining the three basic single-attribute filters using arbitrary logical operators.
Table 2. 2.4) Which filter applies to which attribute.
Unordered attribute Ordered attribute Set attribute
Exact match filter
Range filter
Exact match in set filter
@@ 2.5. Approximate Nearest Neighbor Search
Since many FANNS algorithms build on ANNS methods without filtering, we first introduce the most relevant ANNS algorithms. These can be categorized into tree-based, hash-based, graph-based, and quantization-based methods. For a comprehensive discussion of ANNS, we refer to the surveys by Han et al. (Han et al., 2023) and Echihabi et al. (Echihabi et al., 2021). Experimental evaluations (Aumüller et al., 2020; Aumueller et al., 2025) indicate that no single ANNS algorithm consistently outperforms others; rather, performance depends heavily on dataset characteristics.
@@ 2.5.1. Tree-based methods
Tree-based methods partition the vector space into multiple regions by constructing a search tree, where nodes represent either regions or boundaries between them. Lower tree levels typically correspond to finer partitioning of the vector space. Tree-based methods, such as k-d trees (Bentley, 1975), R-trees (Guttman, 1984), ball trees (Omohundro, 1989), cover trees (Beygelzimer et al., 2006), RP trees (Dasgupta and Freund, 2008), M-trees (Ciaccia et al., 1997), K-means trees (Tavallali et al., 2021), and best bin first (Beis and Lowe, 1997), along with other variants (Houle and Nett, 2014; Lu et al., 2020; Muja and Lowe, 2014; Silpa-Anan and Hartley, 2008), are more efficient for lower-dimensional vectors, as they often suffer from the curse of dimensionality (Marimont and Shapiro, 1979).
@@ 2.5.3. Graph-based methods
Graph-based methods represent vectors as vertices in a graph. Two vertices are connected if their corresponding vectors are close to each other, forming a neighborhood graph (Wang and Li, 2012; Aoyama et al., 2011). For simplicity, we define the length of an edge as the distance between its endpoints’ vectors. To find the nearest neighbors of a query vector, the graph is traversed starting from one or multiple entry points, following the direction of the smallest distance to the query vector. A drawback of neighborhood graphs is that for large datasets, many edges must be traversed, leading to high query times. Navigable small world (NSW) graphs (Malkov et al., 2014) mitigate this issue by introducing long edges that significantly reduce the number of steps needed for traversal. Hierarchical navigable small world (HNSW) graphs (Malkov and Yashunin, 2018) further improve the performance of NSW by introducing a hierarchical structure, where higher layers contain longer edges and lower layers contain shorter ones. The HNSW graph is traversed from the highest to the lowest level. Additional graph-based ANNS methods include DiskANN (Jayaram Subramanya et al., 2019), FreshDiskANN (Singh et al., 2021), NSG (Fu et al., 2017), GRNG (Foster et al., 2025), and others (Zhao et al., 2020; Fu et al., 2021; Harwood and Drummond, 2016; Iwasaki, 2016; Fu et al., 2019).
@@ 2.5.4. Quantization-based methods
[…]
3. Survey and Taxonomy of FANNS
Table 3. 3) Overview of FANNS methods. Does not mention a filtering step during query execution. *We have not been able to find the source code of this work. [Uncaptioned image] The index is constructed for a single ordered attribute, in-filtering is proposed to handle secondary attributes. [Uncaptioned image] Any EMIS filter can be used as an EM filter by using a set attribute with only a single value.
@@ 3. Survey and Taxonomy of FANNS
Rii (Matsui et al., 2018) 2018 FANNS pre-/in-filtering
quantization
(Matsui, 2025)
MA-NSW (Xu et al., 2020) 2019 FANNS hybrid index
graph
PASE (Yang et al., 2020b) 2020 FANNS post-filtering
quantization / graph
(Alipay, 2025)
AnalyticDB-V (Wei et al., 2020) 2020 FANNS pre-/post-filtering
quantization
Milvus (Wang et al., 2021d) 2021 FANNS pre-/post-filtering
quantization / graph
(Project, 2025)
NHQ (Wang et al., 2024, 2022) 2022 AFANNS hybrid index
graph
(Fu, 2025)
HQANN (Yang et al., 2020c) 2022 AFANNS hybrid index
graph
AIRSHIP (Zhao et al., 2022) 2022 FANNS in-filtering
graph
HQI (Mohoney et al., 2023) 2023 FANNS hybrid index
tree + quantization
VBASE (Zhang et al., 2023) 2023 FANNS post-filtering
any
(Microsoft, 2025b)
FDANN (Gollapudi et al., 2023) 2023 FANNS hybrid index
graph
[Uncaptioned image] (Microsoft, 2025a)
CAPS (Gupta et al., 2023) 2023 FANNS hybrid index
quantization + tree
(Gupta, 2025)
ARKGraph (Zuo and Deng, 2023) 2023 FANNS hybrid index
graph
(at Rutgers University, 2025a)
ACORN (Patel et al., 2024) 2024 FANNS in-filtering
graph
(Guestrin Lab at Stanford University, 2025)
$\beta$-WST (Engels et al., 2024) 2024 FANNS hybrid index
tree + graph
(Engels, 2025)
SeRF (Zuo et al., 2024) 2024 FANNS hybrid index
graph
[Uncaptioned image] (at Rutgers University, 2025b)
iRangeGraph (Xu et al., 2024) 2024 FANNS hybrid index
tree + graph
[Uncaptioned image] [Uncaptioned image] (Xu, 2025)
UNIFY (Liang et al., 2024) 2024 FANNS hybrid index
graph
(DBGroup, 2025)
UNG (Cai et al., 2024) 2024 FANNS hybrid index
graph
[Uncaptioned image] [Uncaptioned image] (Cai, 2025)

### v2

@@ Abstract.
Advances in embedding models for text, image, audio, and video drive progress across multiple domains, including retrieval-augmented generation, recommendation systems, and others. Many of these applications require an efficient method to retrieve items that are close to a given query in the embedding space while satisfying a filter condition based on the item’s attributes, a problem known as filtered approximate nearest neighbor search (FANNS). By performing an in-depth literature analysis on FANNS, we identify a key gap in the research landscape: publicly available datasets with embedding vectors from state-of-the-art transformer-based text embedding models that contain abundant real-world attributes covering a broad spectrum of attribute types and value distributions. To fill this gap, we introduce the arxiv-for-fanns dataset of transformer-based embedding vectors for the abstracts of over 2.7 million arXiv papers, enriched with 11 real-world attributes such as authors and categories. We benchmark eleven different FANNS methods on our new dataset to evaluate their performance across different filter types, numbers of retrieved neighbors, dataset scales, and query selectivities. We distill our findings into eight key observations that guide users in selecting the most suitable FANNS method for their specific use cases.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: the 49th International Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne, Australiaisbn: 978-1-4503-XXXX-X/2018/06submissionid: 208
Code: https://github.com/spcl/fanns-benchmark
Dataset: https://hf.co/datasets/SPCL/arxiv-for-fanns-large
@@ 1. Introduction
Advances in embedding models for text (Lee et al., 2024; Lei et al., 2025; Zhang et al., 2024), image (Szegedy et al., 2015; Liu et al., 2024), video (Sivic and Zisserman, 2003a), audio (Cramer et al., 2019; Aytar et al., 2016), and other modalities (Lerer et al., 2019) have significantly enhanced semantic search, where items are mapped to a high-dimensional vector space and similarity is measured by the distance between embedding vectors. Efficient similarity search algorithms are essential to navigate this space. Given the scale of modern datasets, exact nearest neighbor search (ENNS) algorithms are too slow, necessitating the use of approximate nearest neighbor search (ANNS) methods (Indyk and Motwani, 1998; Malkov et al., 2014; Malkov and Yashunin, 2018; Jayaram Subramanya et al., 2019; Sivic and Zisserman, 2003b; Jegou et al., 2010; Gong et al., 2020; Babenko and Lempitsky, 2014a; Guo et al., 2016; Ge et al., 2013). Additionally, many applications, including large language models (LLMs) with retrieval-augmented generation (Gao et al., 2023; Xu et al., 2021; Shapkin et al., 2023), recommendation systems (Suchal and Návrat, 2010; Yang et al., 2020b; Koenigstein et al., 2012), vehicle and person re-identification (Liu et al., 2016a; Zuo and Deng, 2023), and others (Grother et al., 2019; Yang et al., 2020b; Richardson et al., 2015; Li et al., 2018; Wei et al., 2020), require retrieving only items that satisfy filtering conditions on item attributes, such as access rights for a document, a video’s timestamp, or a product’s price. These requirements have driven the development of filtered approximate nearest neighbor search (FANNS). Both academia and industry have recognized its growing importance, leading to numerous research publications (Wang et al., 2024; Mohoney et al., 2023; Gollapudi et al., 2023; Gupta et al., 2023; Zuo and Deng, 2023; Patel et al., 2024; Engels et al., 2024; Zuo et al., 2024; Xu et al., 2024; Liang et al., 2024; Wang et al., 2025; Peng et al., 2025; Luo et al., 2025; Jiang et al., 2025) and adoption in database solutions (Pinecone, 2025; Vespa, 2025; Vectara, 2025).
These efforts require suitable benchmarks to guide the development of FANNS methods, and while transformer-based text embeddings are widely used in practice, no existing FANNS benchmarks with transformer-based embedding vectors exist. Motivated by these developments, we make three main contributions:
First, to clarify the requirements for FANNS benchmarks, we present a comprehensive taxonomy and survey of existing FANNS methods. We identify three key dimensions for the classification of FANNS methods: filtering approach, indexing technique, and supported filter types. To cover all filter types, a FANNS benchmark must contain categorical, numerical, and set-valued attributes, and to reflect the strengths and weaknesses of various filtering approaches, these attributes should exhibit diverse value distributions. The indexing technique is largely orthogonal to the benchmark.
Second, we present the novel arxiv-for-fanns dataset with transformer-based embeddings of paper abstracts from the arXiv dataset (Cornell University et al., 2025) using the model stella_en_400M_v5 (NovaSearch, 2025; Zhang et al., 2024). Our dataset has eleven real-world attributes, including categorical (e.g., venue), numerical (e.g., publication year), and set-valued (e.g., authors) ones, and exhibits diverse value distributions.
Third, we perform a comprehensive benchmarking study of eleven FANNS methods on our novel dataset, evaluating their performance across different filter types, varying numbers of retrieved neighbors, different dataset scales, and a wide range of query selectivities. We distill the results of this benchmarking effort into eight key observations that offer practical guidance for selecting the most appropriate FANNS method for specific application scenarios.
@@ 2. Background
2.1. Filtered Approx. Nearest Neighbor Search
To introduce the FANNS concepts that we leverage throughout the paper, we walk the reader through the process of building and using a semantic search engine for research papers with filtering capabilities, which is a practical use case of FANNS that can be built using our arxiv-for-fanns dataset. Figure 1 provides an overview of this example application, with components and actions referenced using letters [Uncaptioned image] to [Uncaptioned image] and numbers [Uncaptioned image] to [Uncaptioned image], respectively.
Figure 1. 2.1) An Example Application of FANNS.
@@
Index Construction: To build such a system, we transform each of the $n$ papers in our database [Uncaptioned image] into an item. Formally, an item $I_{i}=(v_{i},a_{i,1},\dots,a_{i,m})$ for $i\in\{1,\dots,n\}$ consists of a $d$-dimensional embedding vector $v_{i}$ and a value for each of the $m$ attributes ($a_{i,1}$ to $a_{i,m}$). In our example, the paper’s venue, publication year, and authors define [Uncaptioned image] the item’s attributes. The paper’s abstract is processed [Uncaptioned image] by a text embedding model [Uncaptioned image], such as NV-Embed (Lee et al., 2024; NVIDIA, 2025), LENS (Lei et al., 2025; Lei, 2025), Stella (Zhang et al., 2024; NovaSearch, 2025), GTE (Li et al., 2023; Alibaba-NLP, 2025), or BGE (Chen et al., 2024; of Artificial Intelligence, 2025), to generate the item’s embedding vector $v_{i}$. To efficiently answer FANNS queries for these items, we insert [Uncaptioned image] all items into a FANNS index [Uncaptioned image].
@@
Query execution: During runtime, users interact with the search engine by entering [Uncaptioned image] search terms and filter conditions. We use $p$ to denote the number of user requests submitted during a session. We transform each user request into a FANNS query, formally defined as $Q_{j}=(q_{j},k_{j},f_{j})$ for $j\in\{1,\dots,p\}$ and consisting of a $d$-dimensional query vector $q_{j}$, the number of items to return $k_{j}$, and a filter function $f_{j}$. The filter $f_{j}$ and the number of requested results $k_{j}$ are set [Uncaptioned image] based on the user’s filter conditions. The search term is processed [Uncaptioned image] by the same text embedding model [Uncaptioned image] used during index construction to generate the query vector $q_{j}$. The query is submitted [Uncaptioned image] to a FANNS algorithm [Uncaptioned image], which utilizes [Uncaptioned image] the FANNS index [Uncaptioned image] to retrieve the IDs of the most relevant papers. These IDs are forwarded [Uncaptioned image] to a retriever [Uncaptioned image], which fetches [Uncaptioned image] the corresponding papers from the database [Uncaptioned image] and returns [Uncaptioned image] them to the user.
@@ 2.2. Approximate Nearest Neighbor Search
Since many FANNS algorithms build on ANNS methods without filtering, we first review the most relevant ANNS approaches. They fall into four categories: tree-based, hash-based, graph-based, and quantization-based. For detailed surveys, see Han et al. (Han et al., 2023) and Echihabi et al. (Echihabi et al., 2021). Benchmarks (Aumüller et al., 2020; Aumueller et al., 2025) show that no algorithm dominates across datasets; performance is dataset-dependent.
@@ 2.2.1. Tree-based methods
Tree-based methods partition the vector space into regions using search trees, where nodes denote regions or their boundaries. Deeper levels represent finer partitions. Examples include k-d trees (Bentley, 1975), R-trees (Guttman, 1984), ball trees (Omohundro, 1989), cover trees (Beygelzimer et al., 2006), RP trees (Dasgupta and Freund, 2008), M-trees (Ciaccia et al., 1997), K-means trees (Tavallali et al., 2021), and best bin first (Beis and Lowe, 1997), as well as many variants (Houle and Nett, 2014; Lu et al., 2020; Muja and Lowe, 2014; Silpa-Anan and Hartley, 2008). These methods are effective for low-dimensional vectors but degrade under the curse of dimensionality (Marimont and Shapiro, 1979).
@@ 2.2.3. Graph-based methods
Graph-based methods represent vectors as vertices, connecting those close in the embedding space to form a neighborhood graph (Wang and Li, 2012; Aoyama et al., 2011). Nearest neighbors are found by traversing the graph from one or more entry points, following the shortest distance to the query. Navigable small world (NSW) graphs (Malkov et al., 2014) add long edges between distant vertices to accelerate traversal. Hierarchical navigable small world (HNSW) graphs (Malkov and Yashunin, 2018) extend NSW with a hierarchy: upper layers hold long edges, lower layers short ones, and traversal proceeds top-down. Other graph-based methods include DiskANN (Jayaram Subramanya et al., 2019), FreshDiskANN (Singh et al., 2021), NSG (Fu et al., 2017), GRNG (Foster et al., 2025), and others (Zhao et al., 2020; Fu et al., 2021; Harwood and Drummond, 2016; Iwasaki, 2016; Fu et al., 2019).
@@ 2.2.4. Quantization-based methods
Quantization-based methods map the $n$ vectors to a set of $c\ll n$ clusters, each represented by a cluster centroid. Each vector is assigned to the cluster whose centroid is closest to the vector. One can either store all vectors assigned to a given cluster, as in inverted file (IVF) (Sivic and Zisserman, 2003b), or approximate all vectors of a cluster by the respective cluster centroid, which is less accurate but requires less storage space and allows for faster query times. To find the nearest neighbors of a query vector, we only compare it to the database vectors in the $w\leq c$ clusters whose centroids are closest to the query vector. Product quantization (PQ) (Jegou et al., 2010) and inverted file with product quantization (IVF-PQ) (Jegou et al., 2010) are more advanced quantization-based methods that split the vector space into $s$ orthogonal subspaces and quantize each subspace separately. Other quantization-based ANNS methods include additive quantization (Babenko and Lempitsky, 2014a), BAPQ (Guo et al., 2016), LOPQ (Kalantidis and Avrithis, 2014), OPQ (Ge et al., 2013), RaBitQ (Gao and Long, 2024), SPANN (Chen et al., 2021), and others (Babenko and Lempitsky, 2014b; Wang et al., 2021a; André et al., 2016; Matsui et al., 2015; Wang and Deng, 2020).
Table 1. 3.3) Overview of FANNS methods. Does not mention a filtering step during query execution. *We have not been able to find the source code of this work. [Uncaptioned image] The index is constructed for a single ordered attribute, in-filtering is proposed to handle secondary attributes. [Uncaptioned image] Any EMIS filter can be used as an EM filter by using a set attribute with only a single value.
@@
Rii (Matsui et al., 2018) 2018 FANNS pre-/in-filtering quantization (Matsui, 2025)
MA-NSW (Xu et al., 2020) 2019 FANNS hybrid index graph ✘*
PASE (Yang et al., 2020b) 2020 FANNS post-filtering quantization / graph (Alipay, 2025)
AnalyticDB-V (Wei et al., 2020) 2020 FANNS pre-/post-filtering quantization ✘*
Milvus (Wang et al., 2021d) 2021 FANNS pre-/post-filtering quantization / graph (Project, 2025)
NHQ (Wang et al., 2024, 2022) 2022 AFANNS hybrid index graph (Fu, 2025)
HQANN (Yang et al., 2020c) 2022 AFANNS hybrid index graph ✘*
AIRSHIP (Zhao et al., 2022) 2022 FANNS in-filtering graph ✘*
HQI (Mohoney et al., 2023) 2023 FANNS hybrid index tree + quantization ✘*
VBASE (Zhang et al., 2023) 2023 FANNS post-filtering any (Microsoft, 2025b)
FDANN (Gollapudi et al., 2023) 2023 FANNS hybrid index graph [Uncaptioned image] (Microsoft, 2025a)
CAPS (Gupta et al., 2023) 2023 FANNS hybrid index quantization + tree (Gupta, 2025)
ARKGraph (Zuo and Deng, 2023) 2023 FANNS hybrid index graph (at Rutgers University, 2025a)
ACORN (Patel et al., 2024) 2024 FANNS in-filtering graph (Guestrin Lab at Stanford University, 2025)
$\beta$-WST (Engels et al., 2024) 2024 FANNS hybrid index tree + graph (Engels, 2025)
SeRF (Zuo et al., 2024) 2024 FANNS hybrid index graph [Uncaptioned image] (at Rutgers University, 2025c)
iRangeGraph (Xu et al., 2024) 2024 FANNS hybrid index tree + graph [Uncaptioned image] [Uncaptioned image] (Xu, 2025)
UNIFY (Liang et al., 2024) 2024 FANNS hybrid index graph (DBGroup, 2025)
UNG (Cai et al., 2024) 2024 FANNS hybrid index graph [Uncaptioned image] [Uncaptioned image] (Cai, 2025)
RangePQ (Zhang et al., 2025) 2025 FANNS hybrid index tree + quantization (at The Chinese University of Hong Kong, 2025b)
DIGRA (Jiang et al., 2025) 2025 AFANNS hybrid index tree + graph (at The Chinese University of Hong Kong, 2025a)
DSG (Peng et al., 2025) 2025 FANNS hybrid index graph (at Rutgers University, 2025b)
TFANNS (Luo et al., 2025) 2025 FANNS hybrid index graph [Uncaptioned image] (tfa, 2025)
3. Taxonomy and Survey of FANNS Methods
By analyzing the vast landscape of existing FANNS methods, we identify three key dimensions for classifying these methods. First, they differ in the approach they take to combine ANNS with attribute filtering, which we discuss in Section 3.1. Second, we observe that all FANNS methods are based on the same four fundamental indexing techniques as ANNS methods, namely, hashing, trees, graphs, and quantization, or combinations thereof (see Section 2.2). Third, we distinguish between different filter types that FANNS methods support, which we discuss in Section 3.2. A fourth, less important dimension is the precise problem definition of FANNS that is addressed. Most works guarantee that all returned items pass the filter and only the distance in embedding space is approximated; however, a small number of methods also approximate the filtering step, sometimes returning items that do not match the filter. We refer to this relaxed problem as approximately filtered approximate nearest neighbor search (AFANNS).